Suppose u and v are vectors in 3–space where u = (u1, u2, u3) and v = (v1, v2, v3). Evaluate u×v×u and v × u × u.


1
Expert's answer
2020-09-02T17:44:03-0400

Submit "u=(u_1,u_2,u_3)", "v=(v_1,v_2,v_3)".


Calculate vector product of "u" and "v":

"u\times v=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 u_1 & u_2 & u_3 \\
v_1 & v_2 & v_3
\end{vmatrix}
=\vec i(u_2v_3-u_3v_2)+\vec j(u_1v_3-u_3v_1)+\vec k(u_1v_2-u_2v_1)=\\
{}=(u_2v_3-u_3v_2,\,u_1v_3-u_3v_1,\,u_1v_2-u_2v_1).".


And obtain:

"u\times v\times u=\begin{vmatrix}
 \vec i & \vec j & \vec k \\
 u_2v_3-u_3v_2 & u_1v_3-u_3v_1 & u_1v_2-u_2v_1 \\
 u_1 & u_2 & u_3
\end{vmatrix}
=\vec i(u_1u_3v_3-u_3^2v_1-\\
{}-u_1u_2v_2+u_2^2v_1)+\vec j(u_2u_3v_3-u_3^2v_2-u_1^2v_2+u_1u_2v_1)+\vec k(u_2^2v_3-u_2u_3v_2-\\
{}-u_1^2v_3+u_1u_3v_1)=(u_1u_3v_3-u_3^2v_1-u_1u_2v_2+u_2^2v_1,\,u_2u_3v_3-u_3^2v_2-\\
{}-u_1^2v_2+u_1u_2v_1,\,u_2^2v_3-u_2u_3v_2-u_1^2v_3+u_1u_3v_1)."


According to vector product properties "u\times v\times u=u\times (v\times u)=-(v\times u)\times u", so

"v\times u\times u=-u\times v\times u=-(u_1u_3v_3-u_3^2v_1-u_1u_2v_2+u_2^2v_1,\,u_2u_3v_3-u_3^2v_2-\\
{}-u_1^2v_2+u_1u_2v_1,\,u_2^2v_3-u_2u_3v_2-u_1^2v_3+u_1u_3v_1)=(-u_1u_3v_3+u_3^2v_1+u_1u_2v_2-\\
{}-u_2^2v_1,\,-u_2u_3v_3+u_3^2v_2+u_1^2v_2-u_1u_2v_1,\,-u_2^2v_3+u_2u_3v_2+u_1^2v_3-u_1u_3v_1)."